# Large Language Models Do Not Simulate Human Psychology

### v2

@@ 4.2 Method
Table 1: Morality Ratings for Original and Revised Scenarios by Human Participants and GPT-4
@@
To obtain ratings from LLMs we queried GPT-3.5 (the model used by Dillion et al.), the more recent GPT-4 (in the mini version; one of the LLMs behind ChatGPT), Llama-3.1 70b (the foundation model behind CENTAUR) and CENTAUR with a three-part prompt: First, we provided the same instruction as for human participants, namely: “On the next page, you will be presented with a range of different behaviors. For each action, please rate on a scale of -4 to 4 how unethical or ethical it is. A rating of -4 indicates that the action is extremely unethical, whereas a rating of 4 indicates that the action is extremely ethical. Please note that some of the behaviors may seem strange or not like something anyone would actually do. There might be surprising words or behaviors that seem bizarre. These are not errors. Please still rate how ethical or unethical the action is. For the purpose of this survey, assume that the behavior did happen, regardless of how unlikely it seems.” Second, we provided examples of moral scenarios with ratings (so-called few-shot prompting), in line with the procedure of Dillion et al. (2023). Finally, we provided the new scenarios the model should rate. We repeated each query 10 times to account for random variations and report the mean and standard deviation for humans and GPT-4 (Table LABEL:tab:wordings). GPT-4 is shown here exemplarily as the largest and best-performing LLM among all LLMs included in this paper in terms of the chatbot arena evaluation https://openlm.ai/chatbot-arena/. Data for all ratings by human participants and LLMs can be found at https://osf.io/qbev7/?view_only=1fee15982a714e02bc04ed93cd12afea.
@@ 10. CENTAUR, reworded items -
Table 3: z-Statistics Testing Differences Between Correlation Coefficients. Variables: 1. Human raters, original items, 2. Human raters, reworded items, 3. GPT-3.5, original items, 4. GPT-3.5, reworded items, 5. GPT-4, original items, 6. GPT-4, reworded items, 7. Llama, original items, 8. Llama, reworded items, 9. CENTAUR, original items, 10. CENTAUR, reworded items.
@@ 4.5 LLMs Are Inconsistent in Their Ratings
Figure 3: Rewording-induced change in Morality Ratings. Bar plots show the absolute mean difference between ratings of original scenarios and their reworded counterparts (x-axis), plotted separately for (from top to bottom) human ratings , GPT-3.5, GPT-4,Llama and CENTAUR. To ensure comparability, items are sorted within each panel to reflect the decreasing magnitude of change for human ratings.
Comparing the shifts per scenario in detail (see Figure 3) also reveals inconsistencies between models: GPT4 and GPT-3.5 both tend not to react much to semantic rewording but they react to different scenarios. GPT-3.5 tends to be slightly more in line with human shifts compared to GPT-4. Llama and CENTAUR react more to changes, but quite differently than humans. Changes between Llama and CENTAUR appear consistent, which can be explained by the fact that CENTAUR is a fine-tuned version of Llama.
@@ 4.6 Discussion
Our findings provide a clear demonstration of the limits of using LLMs (at least the four LLMs tested here, GPT-3.5, GPT-4, Llama-3.1 70b and CENTAUR) for simulating human psychology.
First, reproducing the core result of Dillion et al. (2023), we found that LLMs mirror human moral judgments on a set of 30 moral scenarios. This supports the notion that LLMs can replicate human moral judgments on scenarios close to (or contained in) the training data of LLMs. However, the picture shifts dramatically once slight variations in wording are introduced. Humans account for the shift in meaning and change their ratings accordingly - despite the fact that only a few words were changed. By contrast, the ratings of LLMs (especially GPT-3.5 and GPT-4) were hardly affected by the rewordings. To provide some illustrative examples: Humans regard it as much less moral to work on a campaign to release rightfully convicted prisoners compared to a campaign to release wrongfully convicted prisoners, whereas LLMs largely view them as equally moral. Similarly, while human participants viewed setting up traps to catch stray cats as unethical, they viewed it as ethical to set up traps to catch rats. LLMs, on the other hand, viewed both setting traps to catch cats and setting traps to catch rats as unethical. These examples highlight how LLMs can overlook meaningful ethical distinctions that humans make.
@@ 6 Conclusion
The impressive successes of LLMs across a wide range of benchmarks with seemingly human-like or even super-human capabilities has yielded some optimism that LLMs may be able to function as simulators of human cognition, which would equip psychological researchers with a powerful tool to run experiments with large sample sizes without any need for human participants. Perhaps most prominently, the CENTAUR model was recently released with the claim to “predict and simulate human behaviour in any experiment expressible in natural language”. Unfortunately, this is not true (although CENTAUR and Llama-3.1 70b show a higher correlation with human responses in our data compared to GPT-3.5 and GPT-4).

### v3

@@ 4.2 Method
Table 1: Morality Ratings for Original and Revised Scenarios by Human Participants and GPT-4o-mini
@@
To obtain ratings from LLMs we queried GPT-3.5-Turbo (the model used by Dillion et al.), the more recent GPT-4o-mini (one of the LLMs behind ChatGPT), Llama-3.1 70b (the foundation model behind CENTAUR) and CENTAUR with a three-part prompt: First, we provided the same instruction as for human participants, namely: “On the next page, you will be presented with a range of different behaviors. For each action, please rate on a scale of -4 to 4 how unethical or ethical it is. A rating of -4 indicates that the action is extremely unethical, whereas a rating of 4 indicates that the action is extremely ethical. Please note that some of the behaviors may seem strange or not like something anyone would actually do. There might be surprising words or behaviors that seem bizarre. These are not errors. Please still rate how ethical or unethical the action is. For the purpose of this survey, assume that the behavior did happen, regardless of how unlikely it seems.” Second, we provided examples of moral scenarios with ratings (so-called few-shot prompting), in line with the procedure of Dillion et al. (2023). Finally, we provided the new scenarios the model should rate. We repeated each query 10 times to account for random variations and report the mean and standard deviation for humans and GPT-4o-mini (Table LABEL:tab:wordings). GPT-4o-mini is shown here exemplarily as the best-performing LLM among all LLMs included in this paper in terms of the chatbot arena evaluation https://openlm.ai/chatbot-arena/. Data for all ratings by human participants and LLMs can be found at https://osf.io/qbev7/?view_only=1fee15982a714e02bc04ed93cd12afea.
@@ 10. CENTAUR, reworded items -
Table 3: z-Statistics Testing Differences Between Correlation Coefficients. Variables: 1. Human raters, original items, 2. Human raters, reworded items, 3. GPT-3.5, original items, 4. GPT-3.5, reworded items, 5. GPT-4o-mini, original items, 6. GPT-4o-mini, reworded items, 7. Llama, original items, 8. Llama, reworded items, 9. CENTAUR, original items, 10. CENTAUR, reworded items.
@@ 4.5 LLMs Are Inconsistent in Their Ratings
Figure 3: Rewording-induced change in Morality Ratings. Bar plots show the absolute mean difference between ratings of original scenarios and their reworded counterparts (x-axis), plotted separately for (from top to bottom) human ratings , GPT-3.5, GPT-4o-mini,Llama and CENTAUR. To ensure comparability, items are sorted within each panel to reflect the decreasing magnitude of change for human ratings.
Comparing the shifts per scenario in detail (see Figure 3) also reveals inconsistencies between models: GPT4 and GPT-3.5 both tend not to react much to semantic rewording but they react to different scenarios. GPT-3.5 tends to be slightly more in line with human shifts compared to GPT-4o-mini. Llama and CENTAUR react more to changes, but quite differently than humans. Changes between Llama and CENTAUR appear consistent, which can be explained by the fact that CENTAUR is a fine-tuned version of Llama.
@@ 4.6 Discussion
Our findings provide a clear demonstration of the limits of using LLMs (at least the four LLMs tested here, GPT-3.5-Turbo, GPT-4o-mini, Llama-3.1 70b and CENTAUR) for simulating human psychology.
First, reproducing the core result of Dillion et al. (2023), we found that LLMs mirror human moral judgments on a set of 30 moral scenarios. This supports the notion that LLMs can replicate human moral judgments on scenarios close to (or contained in) the training data of LLMs. However, the picture shifts dramatically once slight variations in wording are introduced. Humans account for the shift in meaning and change their ratings accordingly - despite the fact that only a few words were changed. By contrast, the ratings of LLMs (especially GPT-3.5-Turbo and GPT-4o-mini) were hardly affected by the rewordings. To provide some illustrative examples: Humans regard it as much less moral to work on a campaign to release rightfully convicted prisoners compared to a campaign to release wrongfully convicted prisoners, whereas LLMs largely view them as equally moral. Similarly, while human participants viewed setting up traps to catch stray cats as unethical, they viewed it as ethical to set up traps to catch rats. LLMs, on the other hand, viewed both setting traps to catch cats and setting traps to catch rats as unethical. These examples highlight how LLMs can overlook meaningful ethical distinctions that humans make.
@@ 6 Conclusion
The impressive successes of LLMs across a wide range of benchmarks with seemingly human-like or even super-human capabilities has yielded some optimism that LLMs may be able to function as simulators of human cognition, which would equip psychological researchers with a powerful tool to run experiments with large sample sizes without any need for human participants. Perhaps most prominently, the CENTAUR model was recently released with the claim to “predict and simulate human behaviour in any experiment expressible in natural language”. Unfortunately, this is not true (although CENTAUR and Llama-3.1 70b show a higher correlation with human responses in our data compared to GPT-3.5-Turbo and GPT-4o-mini).